# Oversmoothing, “Oversquashing”, Heterophily, Long-Range, and more: Demystifying Common Beliefs in Graph Machine Learning

### v2

@@ 3 Homophily-Heterophily and the Role of the Task
In the context of node classification, the term homophily (resp. heterophily) generally refers to some form of similarity (resp. dissimilarity) between a node and its neighbors [69]. This (dis)similarity can be measured with respect to class labels, node features, or both; the vast majority of works in the literature opt for the first, but this choice is often implicit and taken for granted, making some statements hard to interpret when one is aware of the other ways to measure it.
@@ 3.1 Belief: Homophily is Good, Heterophily is Bad
A recurrent narrative in the literature is that the message-passing mechanism of DGNs is particularly suited for homophilic graphs, whereas it is unfit for heterophilic graphs. The apparent motivation is that, in homophilic graphs, all you need to do to solve a node classification task is to look at similar neighbors, and a local message-passing strategy implements just the right inductive bias. This is in contrast to a class-heterophilic graph, where there exist neighboring nodes of a different class that might make it harder for message-passing to isolate the “relevant information”, intended as neighbors of the same class. Such a belief is supported by empirical evidence on a rather restricted set of benchmarks [86, 73, 77] with varying levels of homophily/heterophily.
@@
Figure 2: Left: a fully heterophilic graph inspired by [67] where a 1-layer, sum-based DGN can perfectly classify the nodes due to a difference in the node degree. Right: a highly homophilic graph where the task is to predict if a node is at a distance greater than five from a specific node. Here, the performances of a DGN will be poor unless information from nodes of another community – from the perspective of a class-0 node – is captured.
Researchers already tried to challenge these considerations in the past [67, 33, 66, 79, 98]. Consider Figure 2 (left), where a bipartite graph with identical node features is fully class-heterophilic. If we apply a single sum-based graph convolution, the nodes can be perfectly classified as the resulting embeddings depend on the incoming degree. Therefore, there exist heterophilic graphs where a DGN can achieve perfect classification. On the contrary, Figure 2 (right) depicts a highly class-homophilic graph, where nodes belong to one of two classes if they are at a distance greater or lower than five from a given node $s$. In this case, the information on the nodes does not even matter; if we were to follow the above belief, we would be encouraged to use a few layers of message-passing, and as a result we could never solve the task perfectly.
@@ 3.2 Belief: Different Classes Imply Different Features
In the previous section, we deliberately ignored the interplay between a node’s features and its class label, which is induced by the task at hand. The reason is that we wanted to clarify the distinction with another, more subtle, and problematic belief: nodes belonging from different classes should have different (to be read as separable) feature distributions. Under this assumption, it is also implicit that class homophily will imply feature homophily.
Such an assumption is often key in arguments supporting the belief of Section 3.1. Indeed, if nodes of different classes have different feature distributions, then applying a local message-passing iteration to a highly class-homophilic graph should “preserve the distance” between node embeddings of different classes. On the contrary, in a heterophilic setting, a graph convolution would mix information coming from different feature distributions, which may be detrimental to performances.
The logic is not incorrect per-se, but our key counterargument is the following: if different classes imply different feature distributions, why would one need to apply a DGN rather than a simple MLP? In other words, either there is a very strong assumption that the task does not depend on the topological information, or the feature-class distributions induced by the task allow us to somehow take a shortcut in terms of learned function, neglecting the role that the topology might have.
It appears therefore necessary to consider less trivial and more fine-grained scenarios, where the feature distributions of different classes partially or totally overlap, the topology has a key role in the task definition, and topological properties induce a positive/negative effect on the performance of message-passing models as done, for instance, in recent works [20, 109].
@@ 3.3 Belief: Long-range Propagation is Evaluated on Heterophilic Graphs
Message of the Section i) Generic claims about the performance of DGNs under homophily and heterophily do not hold, nor does their relation with long-range problems. ii) Homophily/heterophily is a function of the task, but the converse is not true. iii) We should move past the coarse homophily-heterophily dichotomy and focus more on the task and the interplay it induces between features, structure, and class labels.
@@ 4 The Many Facets of “Oversquashing” and their Negative Implications
The term oversquashing originated from Alon and Yahav [4] and referred to an “exponentially growing information into fixed-size vector” by repeated application of message-passing. In other words, oversquashing was associated with the computational tree (Figure 3) induced by message-passing layers on each node of the graph. Later, oversquashing was connected by [93] to the existence of topological bottlenecks: “edges with high negative curvature are those causing the graph bottleneck and thus leading to the over-squashing phenomenon”. Since then, researchers have adopted one or even both definitions of oversquashing at the same time, contributing to an apparent understanding that these definitions subsume the same problem.
@@
Computational bottlenecks, defined in A.1, are inevitably related to the message-passing architecture, for which the graph is the computational medium, whereas topological bottlenecks refer to the graph connectivity. These two problems are clearly intertwined, but in the following we show that they do not always coexist, hence it makes sense to treat them as fundamentally distinct problems.
@@
Figure 3: We intuitively visualize what happens when we repeatedly aggregate the neighborhood of the star node using the message-passing paradigm. Appendix A.4 defines a computational bottleneck as the size of this computational tree (24 computational nodes vs nine 3-hop neighbors).
@@ 4.1 Belief: Oversquashing as Synonym of Topological Bottleneck
The prolific line of work that associates “oversquashing” with topological bottlenecks seems to have gained popularity with Topping et al. [93]. In that paper, edges with negative curvature are first associated with (topological) bottlenecks, then a theorem puts in relation message-passing on a graph containing a bottleneck with the Jacobian sensitivity of node representations, typically defined by $I(u,v)=\lVert\partial\mathbf{h}^{K}_{u}/\partial\mathbf{h}^{0}_{v}\rVert$ and denoting how much the final representation of a node $u$ after $K$ layers is influenced by the initial representation of a node $v$ [102]. Put simply, a topological bottleneck may imply low sensitivity. However, the converse is not necessarily true: we can have low sensitivity on a graph where there are no bottlenecks. Figure 4 (left) shows a grid graph where there are no topological bottlenecks. The repeated application of message passing will, however, quickly generate a computational bottleneck. Therefore, saying that there are no topological bottlenecks does not imply that there are no computational bottlenecks.
To improve on topological bottlenecks, a widely investigated approach is graph rewiring [9], which was also the subject of scrutiny recently [95, 94]. Rewiring is based on the intuition that improving topological bottlenecks metrics should improve the performance of DGNs [6, 12, 52, 28], by reducing the distance between nodes that should communicate. At the same time, it should become clear now that, under the DGN paradigm, rewiring might worsen the computational bottleneck – as long as the same number of message-passing layers is used – while improving the topological one. This perspective was also put forward by Errica et al. [34], with a theoretical analysis on how message filtering, as shown in Figure 2 (middle), reduces both the computational bottleneck and sensitivity yet improves performances while leaving the graph structure unaltered.
@@
Figure 4: Left: in a grid graph, the computational bottleneck grows very quickly but there is no topological bottleneck. Middle: A visualization of the computational graph rooted at node $v_{3}$ for two message passing layers, highlighting how pruning messages reduces the computational bottleneck. Right: In this graph, there is a topological bottleneck and a mild computational bottleneck. As with the grid graph (Appendix B), the sensitivity decreases with the number of message-passing layers.
Another, slightly more technical way to see why low Jacobian sensitivity does not imply the presence of any topological bottlenecks is to follow the chain of upper bounds that link the metrics used to measure the computational and topological bottlenecks [16, 39, 52, 6]. Several recent works [16, 29] have shown that sensitivity is bounded by above by a term that includes the effective resistance, a purely topological distance metric that quantifies the expected commute time of a random walk between nodes $u$ and $v$ [55]. In particular, the bound subtracts the lower bound on the maximum effective resistance, which depends on the inverse of the spectral gap (a proxy for topological bottlenecks) [21, 63]. This connection provides a useful intuition: as the topological bottleneck gets worse, the lower bound on the maximum effective resistance increases, which in turn reduces the upper bounds on the sensitivity of Black et al. [16]. However, the converse does not hold. There are graphs where the maximum effective resistance between distant nodes is large despite the absence of topological bottlenecks. For instance, consider the example of a grid graph in Figure 4 (left) where there is no topological bottleneck, yet the effective resistance between diagonally opposite corners grows linearly with the grid size. As a result, sensitivity between those nodes decays with depth, even though the graph has no identifiable topological bottlenecks. This illustrates that computational bottlenecks can arise independently of topological ones, and that low sensitivity does not necessarily imply the presence of either of them.
@@ 4.2 Belief: Oversquashing as Synonym of Computational Bottleneck
We briefly complement the previous section with a discussion on “oversquashing” as a computational bottleneck, which was introduced by Alon and Yahav [4] and has been the (often implicit) study subject of works that prune, to some extent, the computational tree induced on every node by the iterative message-passing process [81, 34]. Also in this case, there exist cases where reducing the computational bottleneck may be harmful: Figure 4 (right) shows a graph where there is a topological bottleneck but no severe computational bottleneck (for a limited number of layers). In this case, excessive pruning of the computational tree might cause distant nodes to interrupt all communications. Therefore, computational and topological bottlenecks are problems that should be tackled separately.
4.3 Belief: Oversquashing is Problematic for Long-range Tasks
Since its definition by Alon and Yahav [4], oversquashing has often been considered a problem in long-range tasks. The reason stems from its relation to the exponentially growing computational tree as the number of message-passing layers increases: whenever a node has to receive information from another node at distance $d$, classical (synchronous) message-passing architectures need to apply at least $d$ layers to capture that information. As a result, the relevant information
@@ 4.4 Belief: Topological Bottlenecks Associated with Long-range Problems
The last belief we discuss is that topological bottlenecks are the primary obstacle to solving long-range tasks. This intuition stems from the fact that narrow cuts impede information flow between distant parts of the graph. It is indeed true that a topological bottleneck can worsen communication between distant nodes, especially if the bottleneck lies along a path that connects them. However, this perspective is limited in two important ways. First, a topological bottleneck is only harmful if it lies on the information paths between nodes that are supposed to communicate. A topological bottleneck may exist without affecting task-relevant dependencies. Second, the graph topology can worsen the long-range communication even in the absence of any identifiable topological bottleneck, by inducing computational bottlenecks. As we discussed in Section 4.1, the grid graph is an illustrative case: despite the lack of topological bottlenecks, to connect the opposite corner nodes we need, at least, as many message-passing layers as the distance between them, thus leading to a huge computational bottleneck that will likely hamper the ability to process long-range dependencies.
@@
Message of the Section i) “Oversquashing” is an ambiguous term that led to unclear research statements. Talking about computational and topological bottlenecks, instead, better defines the research scope of a paper, since they are two fundamentally distinct problems. ii) There can be computational bottlenecks but no topological ones, and vice-versa. Hence, each of these two bottlenecks, though intertwined, deserves a dedicated research effort. This also means creating ad-hoc benchmarks for each type of bottleneck rather than relying solely on real-world tasks, where it is not as easy to distinguish the combined effect of the two bottlenecks. iii) Computational bottlenecks can happen in short-range as well as long-range tasks. iv) Performance issues in long-range tasks are not solely caused by topological bottlenecks; computational bottlenecks play a role as well.
@@ 5 Conclusions
This position paper posits that the fast pace of advances in the graph machine learning field has generated several commonly accepted beliefs and hypotheses, rooted in the notions of oversmoothing, “oversquashing”, long-range tasks, and heterophily, which are the cause of misunderstandings between researchers. Our contribution was to highlight such beliefs in plain sight, provide an explanation for their emergence, and demystify them when necessary with simple counterexamples. First, we argued that OSM may not be an actual problem and that node embeddings’ separability should be preferred when looking for the root causes of performance degradation. Then, we showed how talking about computational and topological bottlenecks resolves most, if not all, inconsistencies generated by the inflated use of the “oversquashing” term. Finally, we highlighted the role of the task in statements involving homophily, heterophily, and long-range dependencies. By providing much-needed clarifications around these aspects, we hope to foster further advancements in the graph machine learning field.
Appendix A Background
A.1 Deep Graph Networks

### v3

@@ 3 Homophily-Heterophily and the Role of the Task
In the context of node classification, the term homophily (resp. heterophily) generally refers to some form of similarity (resp. dissimilarity) between a node and its neighbors (McPherson et al., 2001). This (dis)similarity can be measured with respect to class labels, node features, or both; the vast majority of works in the literature opt for the first, but this choice is often implicit and taken for granted, making some statements hard to interpret when one is aware of the other ways to measure it.
@@ 3.1 Belief: Homophily is Good, Heterophily is Bad
A recurrent narrative in the literature is that the message-passing mechanism of DGNs is particularly suited for homophilic graphs, whereas it is unfit for heterophilic graphs. The apparent motivation is that, in homophilic graphs, all you need to do to solve a node classification task is to look at similar neighbors, and a local message-passing strategy implements just the right inductive bias. This is in contrast to a class-heterophilic graph, where there exist neighboring nodes of a different class that might make it harder for message-passing to isolate the “relevant information”, intended as neighbors of the same class. Such a belief is supported by empirical evidence on a rather restricted set of benchmarks (Sen et al., 2008; Namata et al., 2012; Pei et al., 2020) with varying levels of homophily/heterophily.
@@
Figure 2: Left: a fully heterophilic graph inspired by Ma et al. (2022) where a 1-layer, sum-based DGN can perfectly classify the nodes due to a difference in the node degree. Right: a highly homophilic graph where the task is to predict if a node is at a distance greater than five from a specific node. Here, the performances of a DGN will be poor unless information from nodes of another community – from the perspective of a class-0 node – is captured.
Researchers already tried to challenge these considerations in the past (Ma et al., 2022; Errica, 2023; Luan et al., 2023; Platonov et al., 2023b; Wang et al., 2024). Consider Figure 2 (left), where a bipartite graph with identical node features is fully class-heterophilic. If we apply a single sum-based graph convolution, the nodes can be perfectly classified as the resulting embeddings depend on the incoming degree. Therefore, there exist heterophilic graphs where a DGN can achieve perfect classification.
On the contrary, Figure 2 (right) depicts a highly class-homophilic graph, where nodes belong to one of two classes if they are at a distance greater or lower than five from a given node $s$. In this case, the information on the nodes does not even matter; if we were to follow the above belief, we would be encouraged to use a few layers of message-passing, and as a result we could never solve the task perfectly. These considerations have motivated the definition of new metrics that put homophily and heterophily in relation with performance under certain assumptions, for instance in Ma et al. (2022).
@@ 3.2 Belief: Different Classes Imply Different Features
In the previous section, we deliberately ignored the interplay between a node’s features and its class label, which is induced by the task at hand. The reason is that we wanted to clarify the distinction with another, more subtle, and problematic belief: nodes belonging to different classes should have different (i.e., separable) feature distributions. Under this assumption, class homophily implies feature homophily.
Such an assumption is often key in arguments supporting the belief of Section 3.1. Indeed, if nodes of different classes have different feature distributions, then applying a local message-passing iteration to a highly class-homophilic graph should “preserve the distance” between node embeddings of different classes. On the contrary, in a heterophilic setting, a graph convolution would mix information coming from different feature distributions, which may be detrimental to performance.
The logic is not incorrect per-se, but our key counterargument is the following: if different classes imply different feature distributions, why would one need to apply a DGN rather than a simple MLP? In other words, either there is a very strong assumption that the task does not depend on the topological information, or the feature-class distributions induced by the task allow us to somehow take a shortcut in terms of learned function, neglecting the role that the topology might have. Note that this discussion is irrespective of the denoising/regularizing effects of DGNs in semi-supervised scenarios compared to MLPs (Hoang et al., 2021; Errica, 2023).
It appears therefore necessary to consider less trivial and more fine-grained scenarios, where the feature distributions of different classes partially or totally overlap, the topology has a key role in the task definition, and topological properties induce a positive/negative effect on the performance of message-passing models as done, for instance, in recent works (Castellana and Errica, 2023; Zheng et al., 2024).
@@ 3.3 Belief: Long-range Propagation is Evaluated on Heterophilic Graphs
Call to Action i) Do not rely on generic claims about the performance of DGNs under homophily and heterophily, nor about their relation with long-range problems. ii) Treat homophily/heterophily properties as task-dependent, not the converse. iii) Move past the coarse homophily-heterophily dichotomy and focus more on the task and the interplay between features, structure, and class labels.
@@ 4 The Many Facets of “Oversquashing” and their Negative Implications
The term oversquashing originated from Alon and Yahav (2021) and referred to an “exponentially growing information into a fixed-size vector” by repeated application of message-passing. In other words, oversquashing was associated with the computational tree (Figure 3) induced by message-passing layers on each node of the graph. Later, oversquashing was connected by Topping et al. (2022) to the existence of topological bottlenecks: “edges with high negative curvature are those causing the graph bottleneck and thus leading to the over-squashing phenomenon”. Since then, researchers have adopted one or even both definitions of oversquashing at the same time, contributing to an apparent understanding that these definitions subsume the same problem.
@@
Computational bottlenecks, defined in Def. 4.1, are inevitably related to the message-passing architecture, for which the graph is the computational medium, whereas topological bottlenecks refer to the graph connectivity, usually measured with spectral or curvature metrics defined in Appendix B.3.2. Therefore, the topological bottlenecks are intrinsic to the graph, while the computational bottleneck is intrinsic to the architecture or procedure. While these bottlenecks are clearly intertwined, in the following we show that they do not always coexist, hence it makes sense to treat them as fundamentally distinct concepts.
Computational and Topological Bottleneck definitions.
Although both bottlenecks have been previously used to measure “oversquashing”, their definitions and measurement approaches also differ across the literature. There is no universally accepted formal definition of a topological bottleneck; instead, it is typically characterized using proxy metrics, most often involving the spectral gap of the curvature. Some of the most significant metrics are summarized in B.3.2.
On the other hand, the computational bottleneck has been usually measured using the receptive field (Chen et al., 2018; Alon and Yahav, 2021), $\mathcal{N}^{k}_{v}$, as the set of $K$-hop neighbors, which grows exponentially with the number of layers.
However, when evaluating the actual computational graph produced by message-passing, shown in Figure 3 (right), duplicate nodes become significant: a node appearing in multiple branches of the computation tree contributes repeatedly, as each unique walk generates a distinct message. Consequently, we employ multiset notation to formally define the computational tree for a node $v$:
Therefore, we can define the notion of an “exponentially-growing receptive field” (Alon and Yahav, 2021) as follows:
Definition 4.1 (Computational Bottleneck).
We refer the reader to B.4 for further insights, related work, and connections between this definition and matrix-power interpretations of message passing.
Figure 3 visualises the difference between the set size $\lvert\mathcal{N}_{v}^{K}\rvert$ and the multiset size $|\mathcal{M}_{v}^{K}|$ on a toy graph and on a simple graph.
@@
Figure 3: We intuitively visualize what happens when we repeatedly aggregate the neighborhood of the star node using the message-passing paradigm, where we define the computational bottleneck as the size of the computational tree (24 computational nodes) for $k=3$ vs the receptive field for $k=3$ that includes 9 three-hop neighbors).
@@ 4.1 Belief: Oversquashing as Synonym of Topological Bottleneck
The prolific line of work that associates “oversquashing” with topological bottlenecks seems to have gained popularity with Topping et al. (2022). In that paper, edges with negative curvature are first associated with (topological) bottlenecks, then a theorem puts in relation message-passing on a graph containing a bottleneck with the Jacobian sensitivity of node representations, typically defined by $I(u,v)=\lVert\partial\mathbf{h}^{K}_{u}/\partial\mathbf{h}^{0}_{v}\rVert$ and denoting how much the final representation of a node $u$ after $K$ layers is influenced by the initial representation of a node $v$ (Xu et al., 2018). Put simply, a topological bottleneck may imply low sensitivity. However, the converse is not necessarily true: we can have low sensitivity on a graph where there are no bottlenecks. Figure 4 (left) shows a grid graph where there are no topological bottlenecks. The repeated application of message passing will, however, quickly generate a computational bottleneck. Therefore, saying that there are no topological bottlenecks does not imply that there are no computational bottlenecks.
To improve on topological bottlenecks, a widely investigated approach is graph rewiring (Attali et al., 2024b), which was also the subject of scrutiny recently (Tortorella and Micheli, 2022; Tori et al., 2025). Rewiring is based on the intuition that improving topological bottlenecks metrics should improve the performance of DGNs (Arnaiz-Rodriguez et al., 2022; Banerjee et al., 2022; Karhadkar et al., 2023; Deac et al., 2022), by reducing the distance between nodes that should communicate. At the same time, it should become clear now that, under the DGN paradigm, rewiring might worsen the computational bottleneck – as long as the same number of message-passing layers is used – while improving the topological one. This perspective was also put forward by Errica et al. (2025), with a theoretical analysis on how message filtering, [^3] The concept of message filtering refers to message passing architectures with the ability to learn how many messages to exchange between nodes and which messages to filter out. as shown in Figure 4 (middle), reduces both the computational bottleneck and sensitivity yet improves performances while leaving the graph structure unaltered.
@@
Figure 4: Left: in a grid graph, the computational bottleneck grows very quickly, but there is no topological bottleneck. Middle: A visualization of the computational graph rooted at node $v_{3}\in\mathcal{V}_{g}$ for two message passing layers, highlighting how pruning messages reduces the computational bottleneck. Right: In this graph, there is a topological bottleneck and a mild computational bottleneck. As with the grid graph (Appendix C), the sensitivity decreases with the number of message-passing layers.
Another, slightly more technical way to see why low Jacobian sensitivity does not imply the presence of any topological bottlenecks is to follow the chain of upper bounds that link the metrics used to measure the computational and topological bottlenecks (Black et al., 2023; Di Giovanni et al., 2023a; Karhadkar et al., 2023; Arnaiz-Rodriguez et al., 2022).
Several recent works (Black et al., 2023; Di Giovanni et al., 2023a) have shown that sensitivity is upper bounded by a term that includes the effective resistance, a purely topological distance metric that quantifies the expected commute time of a random walk between nodes $u$ and $v$ (Klein and Randić, 1993). In particular, the bound subtracts the lower bound on the maximum effective resistance, which depends on the inverse of the spectral gap –a proxy for topological bottlenecks–  (Chandra et al., 1989; Lovász, 1993). This connection provides a useful intuition: as the topological bottleneck gets worse, the lower bound on the maximum effective resistance increases, which in turn reduces the upper bounds on the sensitivity of Black et al. (2023).
However, the converse does not hold. There are graphs where the maximum effective resistance between distant nodes is large despite the absence of topological bottlenecks. For instance, consider the example of a grid graph in Figure 4 (left) where there is no topological bottleneck, yet the effective resistance between diagonally opposite corners grows linearly with the grid size. As a result, sensitivity between those nodes decays with depth, even though the graph has no identifiable topological bottlenecks. This illustrates that computational bottlenecks can arise independently of topological ones, and that low sensitivity does not necessarily imply the presence of either of them.
@@ 4.2 Belief: Oversquashing as Synonym of Computational Bottleneck
We briefly complement the previous section with a discussion on “oversquashing” as a computational bottleneck, which was introduced by Alon and Yahav (2021) and has been the (often implicit) study subject of works that prune, to some extent, the computational tree induced on every node by the iterative message-passing process (Rong et al., 2020; Errica et al., 2025). Also in this case, there exist cases where reducing the computational bottleneck may be harmful: Figure 4 (right) shows a graph where there is a topological bottleneck but no severe computational bottleneck (for a limited number of layers). In this case, excessive pruning of the computational tree might cause distant nodes to interrupt all communications. Therefore, computational and topological bottlenecks are problems that should be tackled separately.
4.3 Belief: Oversquashing is Problematic for Long-range Tasks as well
Since its definition by Alon and Yahav (2021), oversquashing has often been considered a problem in long-range tasks. The reason stems from its relation to the exponentially growing computational tree as the number of message-passing layers increases: whenever a node has to receive information from another node at distance $d$, classical (synchronous) message-passing architectures need to apply at least $d$ layers to capture that information. As a result, the relevant information
@@ 4.4 Belief: Topological Bottlenecks Associated with Long-range Problems
The last belief we discuss is that topological bottlenecks are the primary obstacle to solving long-range tasks. This intuition stems from the fact that narrow cuts impede information flow between distant parts of the graph. It is indeed true that a topological bottleneck can worsen communication between distant nodes, especially if the bottleneck lies along a path that connects them. However, this perspective is limited in two important ways.
First, a topological bottleneck is only harmful if it lies on the information paths between nodes that are supposed to communicate. A topological bottleneck may exist without affecting task-relevant dependencies. Second, the graph topology can worsen the long-range communication even in the absence of any identifiable topological bottleneck, by inducing computational bottlenecks. As we discussed in Section 4.1, the grid graph is an illustrative case: despite the lack of topological bottlenecks, to connect the opposite corner nodes we need, at least, as many message-passing layers as the distance between them, thus leading to a huge computational bottleneck that will likely hamper the ability to process long-range dependencies.
@@
Call to Action i) Stop using the ambiguous “Oversquashing”, which led to unclear research statements. Talking about computational and topological bottlenecks, instead, better defines the research scope of a paper, since they are two fundamentally distinct concepts. ii) There can be computational bottlenecks but no topological ones, and vice-versa. This also means we should create ad-hoc benchmarks for each type of bottleneck rather than relying solely on real-world tasks, where it is not as easy to distinguish the combined effect of the two bottlenecks. iii) Keep in mind that computational bottlenecks can happen in short-range as well as long-range tasks. iv) Do not attribute long-range performance issues solely to topological bottlenecks — account for computational bottlenecks as well.
@@ 5 Conclusions
This paper posits that the fast pace of advances in the graph machine learning field has generated several commonly accepted beliefs and hypotheses, rooted in the notions of oversmoothing, “oversquashing”, long-range tasks, and heterophily, which are the cause of misunderstandings between researchers. Our contribution was to highlight such beliefs in plain sight, provide an explanation for their emergence, and demystify them when necessary with simple counterexamples. First, we argued that OSM may not be an actual problem and that node embeddings’ separability should be preferred when looking for the root causes of performance degradation. Then, we showed how talking about computational and topological bottlenecks resolves most, if not all, inconsistencies generated by the inflated use of the “oversquashing” term. Finally, we highlighted the role of the task in statements involving homophily, heterophily, and long-range dependencies. By providing much-needed clarifications around these aspects, we hope to foster further advancements in the graph machine learning field.
Acknowledgments
A.A. has been partially supported by a nominal grant received at the ELLIS Unit Alicante Foundation from the Regional Government of Valencia in Spain (Resolución de la Conselleria de Industria, Turismo, Innovación y Comercio , Dirección General de Innovación) and a grant by the Banc Sabadell Foundation.
Appendix A Table of Common Beliefs with References
Common Belief and references
Oversmoothing 1. OSM is the cause of performance degradation.
(Li et al., 2018; Hamilton, 2020; Cai and Wang, 2020; Chen et al., 2020b; a; Hasanzadeh et al., 2020; Huang et al., 2020; Liu et al., 2020; Rong et al., 2020; Zhao and Akoglu, 2020; Zhou et al., 2020; Wang and Leskovec, 2021; Zhou et al., 2021; Bodnar et al., 2022; Chen et al., 2022; Hwang et al., 2022; Rusch et al., 2022; Tortorella and Micheli, 2022; Cai et al., 2023; Maskey et al., 2023; Nguyen et al., 2023; Karhadkar et al., 2023; Rusch et al., 2023; ud din and Qureshi, 2024; Wu et al., 2023a; Epping et al., 2024; Jamadandi et al., 2024; Roth and Liebig, 2024; Stanovic et al., 2025; Wang et al., 2025)
2. OSM is a property of all DGNs.
(Xu et al., 2018; Gasteiger et al., 2019; Li et al., 2019; Chen et al., 2020b; Hamilton, 2020; Hasanzadeh et al., 2020; Huang et al., 2020; Zhao and Akoglu, 2020; Zhou et al., 2020; Alon and Yahav, 2021; Zhou et al., 2021; Abboud et al., 2022; Arnaiz-Rodriguez et al., 2022; Chen et al., 2022; Hwang et al., 2022; Keriven, 2022; Topping et al., 2022; Rusch et al., 2022; Akansha, 2023; Cai et al., 2023; Di Giovanni et al., 2023a; Errica et al., 2025; Giraldo et al., 2023; Nguyen et al., 2023; ud din and Qureshi, 2024; Rusch et al., 2023; Shao et al., 2023; Wu et al., 2023a; b; Attali et al., 2024a; b; Fesser and Weber, 2024; Jamadandi et al., 2024; Roth and Liebig, 2024; Arroyo et al., 2025; Wang et al., 2025)
Homophily-Heterophily 3. Homophily is good, heterophily is bad.
(Pei et al., 2020; Zhou et al., 2020; Zhu et al., 2020; Lim et al., 2021; Luan et al., 2021; Wang and Leskovec, 2021; Arnaiz-Rodriguez et al., 2022; Bodnar et al., 2022; Di Giovanni et al., 2023b; Liu et al., 2023b; Platonov et al., 2023a; Bi et al., 2024; Attali et al., 2024a; b; Gong et al., 2024; Roth and Liebig, 2024; Zheng et al., 2024)
4. Long-range propagation is evaluated on heterophilic graphs.
(Arnaiz-Rodriguez et al., 2022; Tortorella and Micheli, 2022; Akansha, 2023; Black et al., 2023; Maskey et al., 2023; Huang et al., 2024; Giraldo et al., 2023; Attali et al., 2024a; Tori et al., 2025)
5. Different classes imply different features.
(Abu-El-Haija et al., 2019; Pei et al., 2020; Lim et al., 2021; Luan et al., 2021; Ma et al., 2022; Sun et al., 2022; Luan et al., 2022; Rusch et al., 2023; Attali et al., 2024a; b; Bi et al., 2024; Huang et al., 2024; Wang et al., 2024; Zheng et al., 2024)
Oversquashing 6. OSQ synonym of a topological bottleneck.
(Xu et al., 2018; Arnaiz-Rodriguez et al., 2022; Banerjee et al., 2022; Chen et al., 2022; Deac et al., 2022; Sun et al., 2022; Topping et al., 2022; Tortorella and Micheli, 2022; Akansha, 2023; Balla, 2023; Black et al., 2023; Di Giovanni et al., 2023a; Errica et al., 2025; Gabrielsson et al., 2023; Giraldo et al., 2023; Karhadkar et al., 2023; Liu et al., 2023a; Nguyen et al., 2023; Shao et al., 2023; Shi et al., 2023; ud din and Qureshi, 2024; Yu et al., 2023; Attali et al., 2024a; b; Barbero et al., 2024; Di Giovanni et al., 2024; Fesser and Weber, 2024; Huang et al., 2024; Jamadandi et al., 2024; Southern et al., 2024; Arroyo et al., 2025; Gravina et al., 2025; Stanovic et al., 2025; Tori et al., 2025)
7. OSQ synonym of computational bottleneck.
(Alon and Yahav, 2021; Abboud et al., 2022; Arnaiz-Rodriguez et al., 2022; Banerjee et al., 2022; Chen et al., 2022; Deac et al., 2022; Sun et al., 2022; Topping et al., 2022; Tortorella and Micheli, 2022; Akansha, 2023; Balla, 2023; Black et al., 2023; Dwivedi et al., 2022; Errica et al., 2025; Giraldo et al., 2023; Gutteridge et al., 2023; Nguyen et al., 2023; Karhadkar et al., 2023; Shao et al., 2023; Shi et al., 2023; ud din and Qureshi, 2024; Attali et al., 2024a; b; Barbero et al., 2024; Huang et al., 2024; Southern et al., 2024; Arroyo et al., 2025; Gravina et al., 2025; Stanovic et al., 2025)
8. OSQ problematic for long-range tasks as well.
(Alon and Yahav, 2021; Abboud et al., 2022; Banerjee et al., 2022; Chen et al., 2022; Deac et al., 2022; Topping et al., 2022; Tortorella and Micheli, 2022; Akansha, 2023; Black et al., 2023; Cai et al., 2023; Di Giovanni et al., 2023a; Errica et al., 2025; Gabrielsson et al., 2023; Karhadkar et al., 2023; Nguyen et al., 2023; Liu et al., 2023a; Shi et al., 2023; Yu et al., 2023; Barbero et al., 2024; Di Giovanni et al., 2024; Fesser and Weber, 2024; Huang et al., 2024; Southern et al., 2024; Attali et al., 2024b; Arroyo et al., 2025; Gravina et al., 2025; Stanovic et al., 2025)
9. Topological bottlenecks associated with long-range problems.
(Topping et al., 2022; Tortorella and Micheli, 2022; Akansha, 2023; Black et al., 2023; Karhadkar et al., 2023; Liu et al., 2023a; Shi et al., 2023; Fesser and Weber, 2024)
Appendix B Background
B.1 Deep Graph Networks